According to a nation wide survey, only 30% of adult Americans had a will. A researcher in Utah suspects that the proportion of adults in Utah with wills is less than the national proportion. Suppose this researcher takes a survey of 1000 adult Utah residents and finds that 277 report having wills. (round your answers to four decimal places)

part b: Calculate and report the value of the test statistic.

part c: Calculate and report the p-value and then use the p-value approach to determine at a 10% level of significance if the proportion of adult Utah residents that have a will is less than the national proportion.

part d: Determine and report the critical value(s) for a 10% level of significance and use the critical value approach to determine if the proportion of adult Utah residents that have a will is less than the national proportion.
Answer #1

Solution :- Let x be the number of American adult had a will in Utah X = 271, n = 1000 Sample proportion ê is. 277 TE 1000 =7 = 1.6812 part co p-value of the test is. P-value = P[z<p cal] =P[< < 1-6312] = 1 -0.0514 p-value = 0.9486 Let Q = 0.10 be t